OpenStudy (anonymous):

A car starts from rest at a stop light and reaches 20 m/s in 3.5 s. Determine the acceleration of the car. A) 3.5 m s2 B) 5.7 m s2 C) 20 m s2 D) 70 m s2

OpenStudy (alexrobin13):

initial velocity: 0 m/s final velocity: 20 m/s time = 3.5s acceleration=(final velocity-initial velocity)/time a = (20-0)/3.5 a = 5.71 m/s^2 So B